OpenStudy (anonymous):

Show that g(x) = -2x - 7/x - 1 = x Please help a girl out!

OpenStudy (31356):

What do you think? o.0

OpenStudy (anonymous):

impossible

OpenStudy (anonymous):

I'm serious..

OpenStudy (anonymous):

\[\frac{-2x-7}{x-1}=x\iff x(x-1)=-2x-7\]\]

OpenStudy (the_fizicx99):

x*x is x^2 ?

OpenStudy (the_fizicx99):

Oh I see what you did! xD

OpenStudy (anonymous):

if it is what i wrote, the solutions are complex (not real numbers)

OpenStudy (anonymous):

I'm slightly confused.

OpenStudy (anonymous):

Aren't there supposed to be a lot of steps?

OpenStudy (anonymous):

this is what you wrote: show \(g(x)=\frac{-2x-7}{x-1}=x\) maybe it is a different question they are not the same thing at all

OpenStudy (anonymous):

can you post a screen shot of the question as written?

OpenStudy (anonymous):

Earlier I was solving for f(g(x)) = x and my steps were: (-2x - 7/x - 1) - 7/(-2x - 7/x - 1) + 2 (-2x - 7/ x - 1) - 7/(-2x - 7/x - 1) = 2 (-2x - 7 -7(x-1)/x - 1)/(-2x - 7 + 2(x -1)/x - 1) -2x - 7 - 7(x-1)/ x - 1 * x - 1/-2x - 7 + 2 (x - 1) -2x - 7 - 7(x-1)/-2x - 7 + 2(x-1) -2x - 7 - 7x + 7/-2x - 7 + 2x - 2 -9x/-9 x

OpenStudy (anonymous):

AHHHH perhaps the question is : let \(g(x)=\frac{-2x-7}{x-1}\) find \(g^{-1}(x)\)?

OpenStudy (anonymous):

\[g(x) = \frac{ -2x - 7 }{ x - 1 } = x\]

OpenStudy (the_fizicx99):

I was also thing that, switching the y for x and x for y ._.

OpenStudy (the_fizicx99):

thinking*

OpenStudy (anonymous):

perhaps it says \[g(x)=\frac{-2x-7}{x-1},f(x)=\frac{x-7}{x+2}\] show \[f(g(x))=x\]???

OpenStudy (anonymous):

It wants me to prove that the equation does indeed equal x

OpenStudy (anonymous):

what equation?

OpenStudy (anonymous):

The g(x) = -2x - 7/ x - 1

OpenStudy (anonymous):

it does not equal \(x\) so you cannot show it

OpenStudy (anonymous):

Hm. Okay. Hang on one sec...

OpenStudy (anonymous):

do you have another function \[f(x)=\frac{x-7}{x+2}\]?

OpenStudy (the_fizicx99):

It'd be better if you screen shot it, g(x) is a function.. >.>

OpenStudy (anonymous):

Yes! I do!

OpenStudy (anonymous):

I should probably just have posted the full question, but I got half done myself so I thought it would be less confusing if I did it this way....here's the question: Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = x - 7/x + 2 and g(x) = -2x - 7/x - 1?

OpenStudy (anonymous):

that is what you did earlier you showed \[f(g(x))=x\] now i think you have to show that \[g(f(x))=x\] does that sound right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

how'd i guess?

OpenStudy (anonymous):

Lol sorry...

OpenStudy (anonymous):

\[g(f(x))\] replace \(f(x)\) by \(\frac{x-7}{x+2}\) and get \[g\left(\frac{x-7}{x+2}\right)\] is the first step

OpenStudy (anonymous):

So I'm writing g(x - 7/x + 2) as the first step...

OpenStudy (anonymous):

then since \[g(\heartsuit)=\frac{-2\heartsuit-7}{\heartsuit-1}\] you have \[g(f(x))=\frac{-2\left(\frac{x-7}{x+2}\right)-7}{\left(\frac{x-7}{x+2}\right)-1}\] as the second step

OpenStudy (anonymous):

What are the hearts supposed to be?

OpenStudy (anonymous):

x's?

OpenStudy (anonymous):

whatever the input is if you use \(x\)'s all the time it gets confusing

OpenStudy (anonymous):

try to think of the function not in terms of \(x\) \[g(x)=\frac{-2x-7}{x-1}\] is the same as \[g(\spadesuit)=\frac{-2\spadesuit-7}{\spadesuit-1}\]

OpenStudy (anonymous):

in any case , now that we have \[g(f(x))=\frac{-2\left(\frac{x-7}{x+2}\right)-7}{\left(\frac{x-7}{x+2}\right)-1}\] it is algebra from here on in

OpenStudy (anonymous):

Okay so so far I have: g(f(x)) = g(x - 7/x + 2) -2 (x - 7/x + 2) - 7/(x - 7/x + 2) - 1 And for now I'm just using x's because it's easier to type....what's next?

OpenStudy (anonymous):

a raft of algebra, then a raft of cancellation

OpenStudy (anonymous):

Could you elaborate?

OpenStudy (anonymous):

\[\frac{-2\left(\frac{x-7}{x+2}\right)-7}{\left(\frac{x-7}{x+2}\right)-1}\] clear the compound fraction by multiplying top and bottom by \(x+2\)

OpenStudy (anonymous):

you get \[\frac{-2(x-7)-7(x+2)}{x-7-1(x+2)}\]

OpenStudy (anonymous):

Okay so I have this so far: g(f(x)) = g(x - 7/x + 2) -2(x - 7/x + 2) - 7/(x - 7/x + 2) - 1 -2(x - 7) - 7(x + 2)/x - 7 - 1(x + 2)

OpenStudy (anonymous):

Then what?

OpenStudy (anonymous):

multiply out in the top and bottom

OpenStudy (anonymous):

\[\frac{-2(x-7)-7(x+2)}{x-7-1(x+2)}\] \[=\frac{-2x+14-7x-14}{x-7-x-2}\]

OpenStudy (anonymous):

then combine like terms top and bottom \[\frac{-9x}{-9}\]

OpenStudy (anonymous):

and cancel, you are done

OpenStudy (anonymous):

So it does equal x? I thought in the beginning you said it doesn't?

OpenStudy (anonymous):

moral of the story post the entire question! it didn't the way you wrote the question at the beginning, without saying to compose the functions and giving the inverse function as \(f(x)=\frac{x-7}{x+2}\)

OpenStudy (anonymous):

Ahhhhh okay. Thank you so much and sorry for the confusion!

OpenStudy (anonymous):

it is not that \[f(x)=\frac{-2x-7}{x-1}=x\] but rather that \[f(g(x))=g(f(x))=x\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Poor satellite lol, I felt your pain as you were trying to figure out the question.

OpenStudy (anonymous):

Can I ask where you are from?